# Cosmological Stasis from Dynamical Scalars: Tracking Solutions and the Possibility of a Stasis-Induced Inflation

### v1

@@ I Introduction, motivation, and overview of results
In this paper, we seek to determine what happens to our stasis phenomenon when we take this full time-dependence into account. At first glance, it might seem that including this time dependence for the equation-of-state parameters for the individual scalar fields would completely destabilize the stasis that emerges when these equation-of-state parameters are instead taken to be constant both before and after the underdamping transition. Indeed, such time-dependent equation of state parameters could in principle complicate the manner in which the Hubble parameter evolves with time, and thus lead to a more complicated dynamical evolution for our scalar fields and their corresponding energy densities. However, since stasis is built on the idea that the abundances of our different fluids remain constant despite cosmological expansion, it is a natural expectation that allowing for time-varying equation-of-state parameters would destroy the stasis that is observed when these equation-of-state parameters are constant.
@@ II A single scalar in an expanding universe
where the effects of the FRW cosmology (i.e., the effects coming from the expansion of the universe) are encoded within the time-dependence of the Hubble parameter $H\equiv\dot{a}/a$. As evident from Eq. (3), the Hubble parameter affects the evolution of the scalar by providing a source of “friction” which damps the motion of the scalar. The value of $H$ is related to the total energy density $\rho_{\rm tot}$ of the universe through the Friedmann equation
@@ II.2 Case II:  Scalar domination
$\phi^{\prime\prime}+\frac{3H}{m}\phi^{\prime}+\phi~{}=~{}0\,$ (9)
@@ III.3 Condition for stasis
The effective equation-of-state parameter $\langle w\rangle$ for the tower as a whole is also essentively constant during stasis. Indeed, this constant value, which we denote ${\overline{w}}$, can be obtained from Eq. (22) by taking $\Omega_{\rm SR}$, $\Omega_{\rm osc}$, and the equation-of-state parameters $w_{\rm SR}$ and $w_{\rm osc}$ equal to their stasis values. In particular, we find that
@@ III.6 Alternative partitions
Thus far we have focused on the partitioning of our tower into only two energy components. However, we can even consider a more general partitioning of the tower into a an arbitrary number $N_{C}$ of cosmological energy components. These components may be labeled by an index $i=1,2,\ldots,N_{C}$. We define an abundance $\Omega_{i}$ and a partition parameter $A_{i}$ for each of these energy components such that $A_{1}=0$ and $A_{i+1}>A_{i}$. For all $i<N_{C}$, we associate the abundances $\Omega_{\ell}$ of all of the $\phi_{\ell}$ with masses within the range $3A_{i+1}H(t)\geq 2m_{\ell}>3A_{i}H(t)$ with $\Omega_{i}$. We associate the abundances $\Omega_{\ell}$ of all the $\phi_{\ell}$ with masses $2m_{\ell}>3A_{N_{C}}H(t)$ with $\Omega_{N_{C}}$.
@@ IV.1 Time-independent background
We begin by investigating the effect that varying both the initial abundance $\Omega_{\rm BG}^{(0)}$ and equation-of-state parameter $w_{\rm BG}$ of the background has on the manner in which the abundances $\Omega_{\rm SR}$ and $\Omega_{\rm osc}$ evolve with time. In Fig. 8, we plot $\Omega_{\rm BG}$ (left panel), $\Omega_{\rm SR}$ (middle panel), and $\Omega_{\rm osc}$ as functions of $m_{0}t$ for several different combinations of $\Omega_{\rm BG}^{(0)}$ and $w_{\rm BG}$.
@@
$\frac{1}{\gamma}\left(\alpha+1/\delta\right)~{}=~{}2-{\overline{\kappa}}~{}=~{}\frac{2\overline{w}}{1+\overline{w}}~{}$ (70)
@@
By contrast, the slow-roll/oscillatory-component stasis that we are discussing in this paper rests upon the much simpler constraint equation in Eq. (39). Indeed, this constraint equation does not involve ${\overline{w}}$ at all, which means that fixing the underlying model parameters $(\alpha,\gamma,\delta)$ does not fix a unique value for ${\overline{w}}$. This in turn means that the properties of the stasis within our dynamical-scalar scenario can be adjusted — even to the extent of changing the value of the equation-of-parameter parameter $\overline{w}$ — while still satisfying our underlying stasis condition. Thus, we see that it is only the slow-roll/oscillatory-component stasis achieved through dynamical scalars which has the freedom needed to “track” a spectator field. This feature thereby distinguishes this stasis from all of the stases that have previously been discussed in the literature. This tracking phenomenon may have important implications when this stasis is embedded in specific cosmological contexts Dienes et al. .
@@ IV.2 Time-dependent backgrounds and tracking solutions
Such tracking behavior is not unexpected. After all, our stasis solution is a global attractor. As a result, any change in $w_{\rm BG}$ simply amounts to moving the location of the attractor within the phase space. Of course, the trajectory lines towards the original attractor are different from the trajectory lines towards the new attractor. However, any point on an original trajectory line is also a point on a new trajectory line. Thus, at the moment when $w_{\rm BG}$ changes, our system simply begins to evolve along the new trajectory line rather the previous one. For this reason, even if the background equation-of-state parameter varies continuously, the tower will always evolve in such a manner as to track $w_{\rm BG}$. Of course, a perfect slow-roll/oscillatory-component stasis with $\langle w\rangle=w_{\rm BG}$ can only be expected once $w_{\rm BG}$ stabilizes to a constant.
@@ V Towards a stasis-induced inflation
Second, we observe that any equation-of-state parameter $\overline{w}<-1/3$ for the stasis sector can be realized within our Stasis Inflation framework. Thus, it is relatively straightforward to achieve an epoch of accelerated cosmological expansion within this scenario, and we are not restricted to having $\overline{w}\approx-1$.
@@ VI Discussion and conclusions
A few comments are in order. First, since the particular form of stasis that we have examined in this paper arises from the collective dynamics of a large number of scalar fields whose masses and abundances exhibit particular scaling behaviors, it is important to consider how these scaling behaviors might arise in actual models of particle physics. Fortunately, the mass spectrum that we considered in this paper — a spectrum characterized by a scaling exponent $\delta$, a mass-splitting parameter $\Delta m$, and a ground-state mass shift $m_{0}$ — is a fairly generic one in many extensions of the Standard Model, including those extensions involving extra compact spacetime dimensions. Moreover, the spectrum of initial abundances that we considered is one in which the $\Omega_{\ell}^{(0)}$ scale with $m_{\ell}$ according to a power law. As we have discussed, this too is a fairly generic result emerging from many different types of production mechanisms.
@@
Second, while we have focused in this paper on the case in which the potentials for the $\phi_{\ell}$ are quadratic, there are also other forms for the potential which are of interest from a model-building perspective. One of these is the form
$V(\phi_{\ell})~{}=~{}\sum_{i}\,\Lambda_{i}\,\exp\left(\sum_{\ell}\alpha_{i\ell}\,\phi_{\ell}\right)~{},$ (79)
@@
where the $\Lambda_{i}$ and $\alpha_{i\ell}$ are model-dependent constants. Potentials of this form arise in the low-energy limit of string compactifications and are of significant interest because they can give rise to so-called scaling cosmologies Calderón-Infante et al. (2023); Shiu et al. (2023a, b) in which the scale factor $a(t)$ evolves with time $t$ according to a power law. While such potentials lack stable minima, it is nevertheless conceivable that they could also give rise to a stasis epoch. Indeed, in such stasis scenarios, the equation-of-state parameter $w_{\ell}$ for each $\phi_{\ell}$ field is approximately $w_{\ell}(t)\approx-1$ at early times. However, since there is no potential minimum, each field continues rolling, and the energy density of the field eventually becomes dominated by kinetic energy. As a result, $w_{\ell}(t)\approx 1$ at late times. It would be interesting to investigate the extent to which this dynamical evolution from smaller to larger values of $w_{\ell}(t)$ can compensate for the effect of cosmological expansion, and whether this effect can therefore also give rise to stasis. We leave this possibility for future work.
@@
Third and finally, we emphasize again that the ideas in this paper could serve as the foundation of a deep connection between cosmology and particle physics and thereby open the door to many exciting phenomenological implications of cosmic stasis. For example, as discussed in Sect. V, a stasis epoch with ${\overline{w}}<-1/3$ persisting for $\mathcal{N}_{s}\gtrsim 60$ $e$-folds of cosmic expansion can in principle constitute a solution to the flatness and horizon problems. If indeed viable models of Stasis Inflation could be developed along these lines, such models would almost certainly give rise to distinctive power spectra of primordial scalar and tensor perturbations. This opens the possibility that evidence for Stasis Inflation could potentially be extracted from observations of the CMB and/or the stochastic gravitational-wave background. Another possibility is that a stasis involving dynamical scalars could occur much later in the cosmological timeline. In particular, it would be interesting to consider the possibility that the slowly rolling $\phi_{\ell}$ could collectively constitute the dark energy which drives the accelerated expansion that we observe at the present time, while the oscillatory $\phi_{\ell}$ could collectively constitute the dark matter. Of course, the presence of large numbers of extremely light axion-like scalars presents a number of model-building challenges, including those imposed by constraints on supernova energy loss Raffelt and Seckel (1988); Olive and Pospelov (2008), Eötvös-type experiments Chen et al. (2016); Tan et al. (2020), searches for frequency variation in atomic clocks Arvanitaki et al. (2015), black-hole superradiance considerations Arvanitaki and Dubovsky (2011); Brito et al. (2015); Cardoso et al. (2018); Stott and Marsh (2018); Stott (2020); Mehta et al. (2020), and data from pulsar-timing arrays Blas et al. (2017); Kaplan et al. (2022). That said, if one were to construct a phenomenologically viable model of dark energy along these lines, it would go a long way toward addressing the cosmic coincidence problem. This topic is therefore worthy of further exploration.

### v2

@@ I Introduction, motivation, and overview of results
In this paper, we seek to determine what happens to our stasis phenomenon when we take this full time-dependence into account. At first glance, it might seem that including this time dependence for the equation-of-state parameters for the individual scalar fields would completely destabilize the stasis that emerges when these equation-of-state parameters are instead taken to be constant both before and after the underdamping transition. Indeed, such time-dependent equation-of-state parameters could in principle complicate the manner in which the Hubble parameter evolves with time, and thus lead to a more complicated dynamical evolution for our scalar fields and their corresponding energy densities. However, since stasis is built on the idea that the abundances of our different fluids remain constant despite cosmological expansion, it is a natural expectation that allowing for time-varying equation-of-state parameters would destroy the stasis that is observed when these equation-of-state parameters are constant.
@@ II A single scalar in an expanding universe
where the effects of the Friedmann-Robertson-Walker cosmology (i.e., the effects coming from the expansion of the universe) are encoded within the time-dependence of the Hubble parameter $H\equiv\dot{a}/a$. As evident from Eq. (3), the Hubble parameter affects the evolution of the scalar by providing a source of “friction” which damps the motion of the scalar. The value of $H$ is related to the total energy density $\rho_{\rm tot}$ of the universe through the Friedmann equation
@@ II.2 Case II:  Scalar domination
$\phi^{\prime\prime}+\frac{3H}{m}\phi^{\prime}+\phi~{}=~{}0\,,$ (9)
@@ III.3 Condition for stasis
The effective equation-of-state parameter $\langle w\rangle$ for the tower as a whole is also essentially constant during stasis. Indeed, this constant value, which we denote ${\overline{w}}$, can be obtained from Eq. (22) by taking $\Omega_{\rm SR}$, $\Omega_{\rm osc}$, and the equation-of-state parameters $w_{\rm SR}$ and $w_{\rm osc}$ equal to their stasis values. In particular, we find that
@@ III.6 Alternative partitions
Thus far we have focused on the partitioning of our tower into only two energy components. However, we can even consider a more general partitioning of the tower into an arbitrary number $N_{C}$ of cosmological energy components. These components may be labeled by an index $i=1,2,\ldots,N_{C}$. We define an abundance $\Omega_{i}$ and a partition parameter $A_{i}$ for each of these energy components such that $A_{1}=0$ and $A_{i+1}>A_{i}$. For all $i<N_{C}$, we associate the abundances $\Omega_{\ell}$ of all of the $\phi_{\ell}$ with masses within the range $3A_{i+1}H(t)\geq 2m_{\ell}>3A_{i}H(t)$ with $\Omega_{i}$. We associate the abundances $\Omega_{\ell}$ of all the $\phi_{\ell}$ with masses $2m_{\ell}>3A_{N_{C}}H(t)$ with $\Omega_{N_{C}}$.
@@ IV.1 Time-independent background
We begin by investigating the effect that varying both the initial abundance $\Omega_{\rm BG}^{(0)}$ and equation-of-state parameter $w_{\rm BG}$ of the background has on the manner in which the abundances $\Omega_{\rm SR}$ and $\Omega_{\rm osc}$ evolve with time. In Fig. 8, we plot $\Omega_{\rm BG}$ (left panel), $\Omega_{\rm SR}$ (middle panel), and $\Omega_{\rm osc}$ (right panel) as functions of $m_{0}t$ for several different combinations of $\Omega_{\rm BG}^{(0)}$ and $w_{\rm BG}$.
@@
$\frac{1}{\gamma}\left(\alpha+1/\delta\right)~{}=~{}2-{\overline{\kappa}}~{}=~{}\frac{2\overline{w}}{1+\overline{w}}~{},$ (70)
@@
By contrast, the slow-roll/oscillatory-component stasis that we are discussing in this paper rests upon the much simpler constraint equation in Eq. (39). Indeed, this constraint equation does not involve ${\overline{w}}$ at all, which means that fixing the underlying model parameters $(\alpha,\gamma,\delta)$ does not fix a unique value for ${\overline{w}}$. This in turn means that the properties of the stasis within our dynamical-scalar scenario can be adjusted — even to the extent of changing the value of the equation-of-state parameter $\overline{w}$ — while still satisfying our underlying stasis condition. Thus, we see that it is only the slow-roll/oscillatory-component stasis achieved through dynamical scalars which has the freedom needed to “track” a spectator field. This feature thereby distinguishes this stasis from all of the stases that have previously been discussed in the literature. This tracking phenomenon may have important implications when this stasis is embedded in specific cosmological contexts Dienes et al. .
@@ IV.2 Time-dependent backgrounds and tracking solutions
Such tracking behavior is not unexpected. After all, our stasis solution is a global attractor. As a result, any change in $w_{\rm BG}$ simply amounts to moving the location of the attractor within the phase space. Of course, the trajectory lines that lead towards the original attractor are different from the trajectory lines that lead towards the new attractor. However, any point on an original trajectory line is also a point on a new trajectory line. Thus, at the moment when $w_{\rm BG}$ changes, our system simply begins to evolve along the new trajectory line rather the previous one. For this reason, even if the background equation-of-state parameter varies continuously, the tower will always evolve in such a manner as to track $w_{\rm BG}$. Of course, a perfect slow-roll/oscillatory-component stasis with $\langle w\rangle=w_{\rm BG}$ can only be expected once $w_{\rm BG}$ stabilizes to a constant.
There is also another potentially interesting aspect of the tracking behavior we have found. Not only does our stasis solution relentlessly track the background when $w_{\rm BG}<\overline{w}$, but it also ceases to track the background if $w_{\rm BG}$ grows too large and exceeds $\overline{w}$. Thus we have at our disposal not only a mechanism for tracking, but also a mechanism for decoupling from tracking! It is therefore possible to arrange things so that our stasis system not only tracks the background for a while but also later decouples from it. This feature can accommodate a broader scope of expansion histories than possible in the traditional scenarios.
@@ V Towards a stasis-induced inflation
Second, we observe that any equation-of-state parameter $\overline{w}<-1/3$ for the stasis sector can be realized within our Stasis Inflation framework. Thus, it is relatively straightforward to achieve an epoch of accelerated cosmological expansion within this scenario, and we are not restricted to having $\overline{w}\approx-1$. Indeed, the ability of our stasis to “track” (and potentially later even decouple from) a time-dependent background may allow the value of $\overline{w}$ to vary throughout the inflationary epoch. This in turn provides us with a handle that might allow us to “dump” unwanted relics (allowing them to inflate away) while simultaneously preserving other relics that might be more desirable.
As an example of this phenomenon, one could imagine inflating away cosmic strings (which have $w=-1/3$) while simultaneously preserving domain or bubble walls (which have $w=-2/3$) so long as $\overline{w}^{\prime}$ lies between these two values. One could then imagine that $\overline{w}^{\prime}$ drops further at a later time, inducing the domain walls to inflate away. In this way we would have not only arranged for both cosmic strings and domain walls to inflate away but also separated the times at which these processes occur within the inflationary epoch. This in turn modifies the manner in which the domain walls associated with bubbles of different sizes exit the horizon, producing results which differ from what is possible in traditional inflation scenarios. Likewise, this also modifies the manner in which a population of bubbles evolves once they subsequently re-enter the horizon. Moreover, in cosmologies of this sort, the changing equation of state of the universe during the inflationary epoch modifies the manner in which density perturbations with different wavenumbers exit the horizon. This can lead to the presence of non-trivial features in the spectrum of density perturbations — features which could leave characteristic imprints in the CMB.
@@ VI Discussion and conclusions
A few comments are in order. First, since the particular form of stasis that we have examined in this paper arises from the collective dynamics of a large number of scalar fields whose masses and abundances exhibit particular scaling behaviors, it is important to consider how these scaling behaviors might arise in actual models of particle physics. Fortunately, the mass spectrum that we considered in this paper — a spectrum characterized by a scaling exponent $\delta$, a mass-splitting parameter $\Delta m$, and a ground-state mass shift $m_{0}$ — is a fairly generic one in many extensions of the Standard Model (SM), including those extensions involving extra compact spacetime dimensions. Moreover, the spectrum of initial abundances that we considered is one in which the $\Omega_{\ell}^{(0)}$ scale with $m_{\ell}$ according to a power law. As we have discussed, this too is a fairly generic result emerging from many different types of production mechanisms.
@@
For simplicity, we have also assumed throughout this paper that while the initial displacements for our $\phi_{\ell}$ fields can be sizable, the corresponding initial velocities $\dot{\phi}_{\ell}^{(0)}$ are sufficiently small that their impact on the subsequent evolution of the system can be neglected. There are a number of ways in which such initial conditions for the $\dot{\phi}_{\ell}^{(0)}$ can be achieved. One possibility, which is discussed in Ref. Dienes et al. (2000), arises within the context of theories with extra spacetime dimensions in which the SM is localized on a four-dimensional brane, while an additional scalar field neutral under the SM gauge group propagates freely within the higher-dimensional bulk. The KK modes of this bulk scalar play the role of the $\phi_{\ell}$ in this context. At early times, the mass eigenstates of the theory are simply the KK-number eigenstates. One natural set of initial conditions for the zero-modes of these scalars, which follows from positing that by some early time they have all effectively settled into their potential minima, is $\phi_{\ell}\approx 0$ and $\dot{\phi}_{\ell}\approx 0$ for all of the $\phi_{\ell}$ except $\phi_{0}$, which is massless and can therefore acquire a misaligned background value. If additional terms in the mass matrix for the $\phi_{\ell}$ are generated dynamically at some later time — for example, as the consequence of a second-order phase transition — the mass eigenstates of the theory are subsequently admixtures of the KK-number eigenstates. These late-time eigenstates acquire non-zero field displacements (but negligible field velocities) via their mixing with the KK zero-mode.
Another possibility is that the $\dot{\phi}_{\ell}$ are initially non-negligible at very early times. However, the kinetic-energy density $\rho_{\ell}^{({\rm KE})}$ associated with each of these field velocities damps away rapidly in comparison with the corresponding potential energy. As a result, the universe will ultimately tend toward a state in which all of the $\phi_{\ell}$ are either slowly rolling or oscillating around their potential minima. This suppression of kinetic-energy density occurs generically, regardless of the effective equation-of-state for the universe prior to the onset of stasis, though the rate at which the $\rho_{\ell}^{({\rm KE})}$ are suppressed of course depends on this equation-of-state parameter. One possibility which would result in a particularly efficient suppression of the $\rho_{\ell}^{({\rm KE})}$ involves positing that a brief auxiliary period of accelerated cosmic expansion took place before stasis began. Such a period of accelerated expansion could potentially even be driven by the $\phi_{\ell}$ themselves, provided that $H^{(0)}/m_{N-1}$ is sufficiently large. During this period, the $\rho_{\ell}^{({\rm KE})}$ would decrease exponentially with $t$. We emphasize this period need not last more than a few $e$-folds in order to establish an appropriate set of initial conditions for the $\phi_{\ell}$, and therefore would not usurp the role of the subsequent stasis epoch in providing a solution to the horizon problem.
Second, spatially homogeneous scalar-field configurations are generically unstable due to parametric resonances associated with self-interaction terms in the scalar potential (if such terms are present) Kofman et al. (1994); Shtanov et al. (1995); Kofman et al. (1997) or to gravitational instabilities Jedamzik et al. (2010); Easther et al. (2011); Musoke et al. (2020). As a result, the homogeneous zero-mode configurations of our $\phi_{\ell}$ are expected to fragment over time into a collection of spatially localized, gravitationally bound structures or “lumps.” However, the fragmentation of the initially homogeneous field configurations of our $\phi_{\ell}$ should not have a significant impact on the dynamics which gives rise to stasis. Indeed, this fragmentation, which for the purely quadratic form of $V_{\ell}$ we are considering in this paper is due to gravitational instabilities, occurs only after the field in question begins oscillating Jedamzik et al. (2010). Moreover, for this form of $V_{\ell}$, the lumps into which each $\phi_{\ell}$ condensate fragments behave like massive matter. Thus, both before and after fragmentation, each $\phi_{\ell}$ which satisfies $3H(t)<2m_{\ell}$ at a given time $t$ and therefore contributes to $\Omega_{\rm osc}$ has an equation-state parameter $w_{\ell}\approx 0$. Fragmentation therefore has no fundamental impact on the cosmological dynamics which gives rise to stasis in this case, irrespective of the value of ${\overline{w}}$. This conclusion still holds even if the $V_{\ell}$ potentials are not purely quadratic, but include additional terms which are higher-order in $\phi_{\ell}$, provided that the anharmonicities are sufficiently small.
For other forms of $V_{\ell}$ (e.g., for quartic potentials), the gravitationally bound structures into which the $\phi_{\ell}$ condensates fragment do not necessarily behave like massive matter. That said, in the inflationary regime in which ${\overline{w}}\leq-1/3$, the energy density associated with these structures rapidly redshifts away during stasis. Since each individual field only contributes significantly to $\Omega_{\rm osc}$ for a brief period after it has begun oscillating, and since fragmentation of each $\phi_{\ell}$ condensate in this case likewise occurs only once the field has begun oscillating, the impact of fragmentation on a stasis with such a value of ${\overline{w}}$ may also be negligible in many cases, depending on the rate at which $\phi_{\ell}$ particles are produced via parametric resonance. We leave a detailed study of this dynamics for future work.
@@
Third, while we have focused in this paper on the case in which the potentials for the $\phi_{\ell}$ are quadratic, there are also other forms for the potential which are of interest from a model-building perspective. One of these is the form
$V(\phi_{\ell})~{}=~{}\sum_{i}\,\Lambda_{i}\,\exp\left(\sum_{\ell}\alpha_{i\ell}\,\phi_{\ell}\right)~{},~{}~{}$ (79)
@@
where the $\Lambda_{i}$ and $\alpha_{i\ell}$ are model-dependent constants. Potentials of this form arise in the low-energy limits of various string compactifications and are of significant interest because they can give rise to so-called scaling cosmologies Shiu et al. (2023a, b) (see also Ref. Calderón-Infante et al. (2023)) in which the scale factor $a(t)$ evolves with time $t$ according to a power law. While such potentials lack stable minima, it is nevertheless conceivable that they could also give rise to a stasis epoch. Indeed, even in scenarios involving a single scalar field with an exponential potential Copeland et al. (1998), one finds that the equation-of-state parameter $w$ for the single scalar field $\phi$ is approximately $w(t)\approx-1$ at early times. However, since there is no potential minimum, this field gathers kinetic energy as it rolls, and as a result $w(t)$ increases. It would be interesting to investigate how this behavior is modified in a multi-field context and to examine the extent to which the dynamical evolution from smaller to larger values of $w_{\ell}(t)$ can compensate for the effect of cosmological expansion, thereby potentially giving rise to a stasis epoch. We leave this possibility for future work.
@@
Fourth and finally, we emphasize again that the ideas in this paper could serve as the foundation of a deep connection between cosmology and particle physics and thereby open the door to many exciting phenomenological implications of cosmic stasis. For example, as discussed in Sect. V, a stasis epoch with ${\overline{w}}<-1/3$ persisting for $\mathcal{N}_{s}\gtrsim 60$ $e$-folds of cosmic expansion can in principle constitute a solution to the flatness and horizon problems. If indeed viable models of Stasis Inflation could be developed along these lines, such models would almost certainly give rise to distinctive power spectra of primordial scalar and tensor perturbations. This opens the possibility that evidence for Stasis Inflation could potentially be extracted from observations of the CMB and/or the stochastic gravitational-wave background. Another possibility is that a stasis involving dynamical scalars could occur much later in the cosmological timeline. In particular, it would be interesting to consider the possibility that the slowly rolling $\phi_{\ell}$ could collectively constitute the dark energy which drives the accelerated expansion that we observe at the present time, while the oscillatory $\phi_{\ell}$ could collectively constitute the dark matter. Of course, the presence of large numbers of extremely light axion-like scalars presents a number of model-building challenges, including those imposed by constraints on supernova energy loss Raffelt and Seckel (1988); Olive and Pospelov (2008), Eötvös-type experiments Chen et al. (2016); Tan et al. (2020), searches for frequency variation in atomic clocks Arvanitaki et al. (2015), black-hole superradiance considerations Arvanitaki and Dubovsky (2011); Brito et al. (2015); Cardoso et al. (2018); Stott and Marsh (2018); Stott (2020); Mehta et al. (2020), data from pulsar-timing arrays Blas et al. (2017); Kaplan et al. (2022), and the inverse correlation which exists between the dark-energy equation of state and the present-day value of $H$ in combined fits to CMB and supernova data Zhao et al. (2017); Vagnozzi et al. (2018); Vagnozzi (2020); Banerjee et al. (2021); Lee et al. (2022). That said, if one were to construct a phenomenologically viable model of dark energy along these lines, it would go a long way toward addressing the cosmic coincidence problem. This topic is therefore worthy of further exploration.